A cheetah is running at a speed of 20.4 m/s in a d
Answer #1

20.4 m/s v.sin790 790 v.cos790

Part A) North Component of the velocity

v_{N}= (20.4 m/s)\sin 79^{0} = 20.025 m/s

Part B) West Component of the velocity

v_{W}= (20.4 m/s)\cos 79^{0} = 3.8925 m/s